# CodeARC: Benchmarking Reasoning Capabilities of LLM Agents for Inductive Program Synthesis

### v1

@@ Abstract
Inductive program synthesis, or programming by example, requires synthesizing functions from input-output examples that generalize to unseen inputs. While large language model agents have shown promise in programming tasks guided by natural language, their ability to perform inductive program synthesis is underexplored. Existing evaluation protocols rely on static sets of examples and held-out tests, offering no feedback when synthesized functions are incorrect and failing to reflect real-world scenarios such as reverse engineering. We propose CodeARC, the Code Abstraction and Reasoning Challenge, a new evaluation framework where agents interact with a hidden target function by querying it with new inputs, synthesizing candidate functions, and iteratively refining their solutions using a differential testing oracle. This interactive setting encourages agents to perform function calls and self-correction based on feedback. We construct the first large-scale benchmark for general-purpose inductive program synthesis, featuring 1114 functions. Among 18 models evaluated, o3-mini performs best with a success rate of 52.7%, highlighting the difficulty of this task. Fine-tuning LLaMA-3.1-8B-Instruct on curated synthesis traces yields up to a 31% relative performance gain. CodeARC provides a more realistic and challenging testbed for evaluating LLM-based program synthesis and inductive reasoning.
@@ 1 Introduction
Inductive reasoning, i.e., the ability to identify patterns and form abstractions from limited examples, is widely recognized as a fundamental aspect of human intelligence (Hayes et al., 2010; Yan et al., 2025). In the context of programming, inductive reasoning underpins the task of synthesizing functions that satisfy given input-output examples and generalize to unseen inputs. This task, commonly referred to as inductive program synthesis or programming by example (Gulwani, 2011; Feser et al., 2023; Li & Ellis, 2025), has broad application domains (Wang et al., 2017; Deng et al., 2024).
@@
Designing an effective evaluation protocol for inductive program synthesis with LLMs is inherently challenging, as multiple valid functions may satisfy a given set of input-output examples (as demonstrated in Section 5.2). The state-of-the-art protocol (Li & Ellis, 2025), which evaluates synthesized functions on held-out test cases after presenting 10 fixed input-output examples, has several notable limitations. First, a small, static set of input-output examples may underspecify the target functions, especially for those with complex logic. Moreover, held-out tests may fail to reveal subtle semantic discrepancies between the generated and intended implementations. In addition, when the model produces an incorrect solution, it receives no feedback and has no opportunity to revise or explore alternatives. Finally, existing benchmarks for inductive program synthesis (Gulwani, 2011; Wang et al., 2017; Wong et al., 2021; Deng et al., 2024; Li & Ellis, 2025) are narrowly focused on domain-specific tasks and do not assess the ability of LLMs to synthesize functions written in general-purpose programming languages.
@@
To address the limitations of existing evaluation protocols for inductive program synthesis, we introduce CodeARC, the Code Abstraction and Reasoning Challenge, inspired by real-world scenarios such as decompilation and reverse engineering (Lin et al., 2010; Mantovani et al., 2022; Al-Kaswan et al., 2023). In such settings, an agent is given a binary executable (without source code) and must synthesize an equivalent source code by observing input-output behavior. Rather than relying on a fixed dataset, the agent can query the binary with new inputs, invoke a differential testing oracle to compare the synthesized program against the reference binary, and use counterexamples to iteratively refine its solution.
As shown in Figure 1, our framework CodeARC emulates this process: LLM-based agents interact with an environment by starting with an initial set of input-output examples, querying the ground-truth function for additional examples, and debugging synthesized code using a differential testing oracle. We impose fixed budgets on both the number of observable input-output examples and the number of testing oracle invocations for self-debugging. The task requires agents to proactively generate inputs (function calls) and revise solutions based on feedback (self-reflection). This interactive setup offers a more realistic alternative to prior static evaluation protocols.
We construct the first comprehensive dataset for general-purpose inductive program synthesis, featuring 1114 functions with initial input-output examples. It targets general programming tasks and employs two state-of-the-art differential testing tools (Lukasczyk & Fraser, 2022; Etemadi et al., 2024) for correctness evaluation.
@@ 2 Related Work
Most LLM benchmarks—such as HumanEval+ (Chen et al., 2021; Liu et al., 2023b), MBPP+ (Austin et al., 2021; Liu et al., 2023b), APPS (Hendrycks et al., 2021), and others (Li et al., 2022; Liu et al., 2023c; Li et al., 2023; Jain et al., 2024a; Banerjee et al., 2025; Ni et al., 2024b; Liu et al., 2023d; Du et al., 2023; Suresh et al., 2025b; Zhuo et al., 2024; Yin et al., 2022; Lai et al., 2023; Patil et al., 2025; Ugare et al., 2024; 2025)—evaluate code generation from natural language. Beyond generation, tasks like I/O prediction (Gu et al., 2024; Liu et al., 2024a), execution prediction (Liu et al., 2023a; La Malfa et al., 2024; Ni et al., 2024a; Ding et al., 2024), bug localization (Suresh et al., 2025a), and program equivalence (Wei et al., 2025) have also been studied. In contrast, we focus on predicting function bodies purely from input-output examples, without natural language. Prior work (Li & Ellis, 2025; Barke et al., 2024) targets domain-specific tasks, while we introduce a general-purpose benchmark with an interactive evaluation protocol.
@@
LLMs are widely benchmarked on reasoning tasks across domains, including commonsense (Talmor et al., 2018), mathematical (Cobbe et al., 2021), and logical (Han et al., 2022; Miao et al., 2021; Liu et al., 2020; Parmar et al., 2024). Inductive reasoning, a core cognitive skill that generalizes from limited examples (Hayes et al., 2010), is increasingly studied in LLMs (Li et al., 2024; Ma et al., 2024; Xiao et al., 2024; Cai et al., 2024). ARC (Chollet, 2019) is a prominent benchmark for abstract pattern induction. Our work shares this goal but is for inductive program synthesis.
@@
LLM-based agents have shown strong performance in domains like web navigation (Zhou et al., 2024a; b), code generation (Zhang et al., 2023; Jimenez et al., 2024), and ML experimentation (Huang et al., 2024). They interact with environments, invoke functions, make decisions, and self-reflect (Madaan et al., 2023; Paul et al., 2023; Xie et al., 2023; Huang et al., 2022; Shinn et al., 2023). We introduce the first benchmark to systematically evaluate agents’ capabilities in inductive program synthesis, providing a rigorous testbed for inductive reasoning and program synthesis.
@@ 3.1 Problem Definition of Inductive Program Synthesis
To evaluate whether a synthesized function $\hat{f}$ is correct, we introduce a differential testing oracle $\mathcal{O}$. The oracle takes as input both the synthesized function $\hat{f}$ and the hidden ground-truth function $f^{*}$ and attempts to identify inputs on which their behaviors diverge. Formally, the oracle operates as follows:
@@
Unlike fixed held-out test sets used in prior work, the differential testing oracle conditions on both $f^{*}$ and the candidate $\hat{f}$, generating targeted inputs to expose discrepancies. On failure, it returns a counterexample $x\in\mathcal{X}_{\text{test}}$ such that $\hat{f}(x)\neq f^{*}(x)$. Note that program equivalence checking is fundamentally undecidable, and thus no perfect oracle exists. To approximate oracle functionality, we adopt two state-of-the-art differential testing tools, providing a more robust and practical evaluation.
@@ 3.2 Interactive Evaluation Protocol for LLM Agents
During evaluation, the agent may take two types of actions. First, it may query the ground-truth function $f^{*}$ at a chosen input $x\in\mathcal{X}$, and obtain the corresponding output $f^{*}(x)$, thereby augmenting its observed set of input-output pairs. Second, it may synthesize a candidate program $\hat{f}$ and invoke the differential testing oracle $\mathcal{O}(f^{*},\hat{f})$, which returns Pass if no discrepancies are found on a dynamically generated test set, or Fail with a counterexample $x\in\mathcal{X}_{\text{test}}$ such that $\hat{f}(x)\neq f^{*}(x)$.
@@
The task is considered successful if the final synthesized program $\hat{f}$, produced within the given budgets, receives a Pass from the differential testing oracle $\mathcal{O}(f^{*},\hat{f})$. We assess the performance of an LLM agent along two dimensions: correctness and efficiency, with correctness being the primary criterion. The objective is to synthesize correct programs, even at the cost of increased resource consumption. Correctness is quantified by the success rate, defined as the proportion of problems in the dataset for which the agent successfully synthesizes a correct program within the specified budget. Efficiency is measured by the average number of input-output queries and oracle invocations utilized per problem.
@@ 3.3 Benchmark Preparation
We generate a synthetic dataset for fine-tuning (described in more detail in Section 3.4) by first collecting 50 seed Python functions that are disjoint from the evaluation set. Using these seeds, we prompt GPT-4o to synthesize a diverse set of new functions, yielding 10,000 candidates. For each generated function, we additionally instruct GPT-4o to produce 10 representative inputs that expose the function’s behavior and highlight patterns in its input-output relationships. These inputs are executed to verify the executability of the functions, and we discard any that fail at this stage. To reduce redundancy, we deduplicate the dataset based on function names, finally resulting in 5,405 unique Python functions used for fine-tuning.
@@ 3.4 Fine-Tuning on Synthetic Data
We investigate whether simply fine-tuning on curated synthesis traces, which capture both function invocations and reasoning steps, can improve LLM performance on CodeARC. We adopt a knowledge distillation fine-tuning approach, where the student model learns to imitate the reasoning and behavior of a teacher model that has access to the ground-truth function body of $f^{*}$.
During training, we first run the interactive evaluation protocol described in Section 3.2 with a frozen teacher model. Unlike a standard evaluation, we prepend a set of task-specific instructions $P_{f^{*}}$ to each teacher prompt. This prefix includes the function body of $f^{*}$ and explicitly instructs the teacher to (1) query $f^{*}$ on informative inputs, (2) explain the rationale behind those queries, and (3) synthesize the full implementation of $f^{*}$ only when confident that the correct logic can be inferred from the accumulated input-output pairs. The student model, which is the only model we fine-tune, learns to mimic the teacher’s reasoning and synthesis behavior without ever seeing the function body.
@@ 4 Experiment Setup
We construct two versions (annotated and anonymized) of the 1114 Python functions, drawn from HumanEval+, MBPP+ (Liu et al., 2023b), and APPS (Hendrycks et al., 2021). Table 1 summarizes key statistics. Unlike prior work in program synthesis that often focuses on domain specific languages and constrained settings, our benchmark consists of programs written in Python, a general purpose language that captures a broader range of real-world algorithms and tasks.
For the main evaluation (Section 5.1), we provide 10 initial input-output examples and set the query budget to 30 input-output pairs and 2 oracle calls ($B_{\text{io}}=30$, $B_{\text{oracle}}=2$), chosen based on practical constraints such as API cost and runtime. Section 5.3 reports ablation studies on both budgets. We use two state-of-the-art differential testing tools, Pynguin (Lukasczyk & Fraser, 2022) and Mokav (Etemadi et al., 2024). For supervised fine-tuning on synthetic reasoning trajectories, we use gpt-4o as the teacher model and LLaMA-3.1-8B-Instruct as the student. See Appendix A for further experiment details, including prompts and the fine-tuning parameters.
@@ 5.1 Main Results
Table 2 shows the results for 18 large language models on CodeARC. The order is sorted based on the success rate on the anonymized dataset. We also report the average number of observed input-output examples and oracle invocations. Our findings are as follows:
@@
Larger models generally achieve better performance, as shown in Figure 2 (log-scale x-axis). All model families exhibit scaling trends, though with varying consistency. Llama-3.1 scales steadily, while Llama-3.2 plateaus at larger sizes, likely due to its multimodal focus. Qwen2.5 shows clearer scaling on anonymized data, where reasoning is required over memorization, highlighting model size’s impact on generalization.
@@ 5.2 Do Initial Input-Output Examples Underspecify the Target Function?
To assess whether 10 input-output examples (Li & Ellis, 2025) suffice to specify the target function, we evaluate the first synthesized function of o3-mini—the strongest model in our study. Functions that pass the initial examples but fail under oracle testing indicate under-specification, motivating the need for additional examples or oracle-guided feedback. As shown in Table 4, 67.6% pass the initial 10 input-output example tests, but only 38.9% pass the oracle, revealing 640 cases (28.7%) where the initial examples fail to uniquely specify the target function. These findings show that initial input-output examples often under-specify program behavior, motivating additional queries and oracle-guided feedback for reliable evaluation. This motivates the design of our interactive evaluation protocol.
@@ 5.3 Ablation Study: Input-Output Queries and Oracle Feedback
We evaluate o3-mini and o1-mini with input-output budgets of 10, 20, and 30, using the same setup as Section 5.1. Table 4 shows that success rates improve consistently with more examples.
@@
We vary the number of allowed oracle invocations and report success rates in Figure 3 for four models on both datasets. More oracle calls consistently improve performance, showing that counterexamples from differential testing are valuable for guiding iterative refinement.
@@ 5.4 Performance of Fine-Tuned Models
Table 5 shows that fine-tuning the LLaMA-3.1-8B-Instruct model on curated synthesis traces yields consistent improvements across both datasets. The larger gain on the annotated variant suggests that fine-tuning is particularly effective when semantically informative identifiers are present. Notably, this performance gap emerges despite both datasets being evaluated under the same model architecture and training methodology discussed in Section 3.4. These results indicate that while fine-tuning helps, there remains substantial room for further improvement. This suggests that future research may focus on enhancing the quality and diversity of the fine-tuning dataset, particularly for the anonymized variant where gains are more limited. Another promising direction is to explore reinforcement learning approaches that directly optimize for higher synthesis success rates, potentially overcoming limitations of supervised fine-tuning alone.

### v2

@@ Abstract
Inductive program synthesis, or programming by example, requires synthesizing functions from input-output examples that generalize to unseen inputs. While large language model agents have shown promise in programming tasks guided by natural language, their ability to perform inductive program synthesis is underexplored. Existing evaluation protocols rely on static sets of examples and held-out tests, offering no feedback when synthesized functions are incorrect and failing to reflect real-world scenarios such as reverse engineering. We propose CodeARC, the Code Abstraction and Reasoning Challenge, a new evaluation framework where agents interact with a hidden target function by querying it with new inputs, synthesizing candidate functions, and iteratively refining their solutions using a differential testing oracle. This interactive setting encourages agents to perform function calls and self-correction based on feedback. We construct the first large-scale benchmark for general-purpose inductive program synthesis, featuring 1114 functions. Among 18 models evaluated, o3-mini performs best with a success rate of 52.7%, highlighting the difficulty of this task. Fine-tuning LLaMA-3.1-8B-Instruct on curated synthesis traces yields up to a 31% relative performance gain. CodeARC provides a more realistic and challenging testbed for evaluating LLM-based program synthesis and inductive reasoning. Our code, data, and models are publicly available at https://github.com/Anjiang-Wei/CodeARC
footnotetext: Correspondence to: anjiang@cs.stanford.edu
@@ 1 Introduction
Inductive reasoning, i.e., the ability to identify patterns and form abstractions from limited examples, is widely recognized as a fundamental aspect of human intelligence (Hayes et al., 2010; Yan et al., 2025). In the context of programming, inductive reasoning underpins the task of synthesizing functions that satisfy given input-output examples and generalize to unseen inputs. This task, commonly referred to as inductive program synthesis or programming by example (Manna & Waldinger, 1971; Myers, 1986; Feser et al., 2023; Li & Ellis, 2025), has broad application domains (Wang et al., 2017; Deng et al., 2024).
@@
Designing an effective evaluation protocol for inductive program synthesis with LLMs is inherently challenging, as multiple valid functions may satisfy a given set of input-output examples (as demonstrated in Section˜5.2). The state-of-the-art protocol (Li & Ellis, 2025), which evaluates synthesized functions on held-out test cases after presenting 10 fixed input-output examples, has several notable limitations. First, a small, static set of input-output examples may underspecify the target functions, especially for those with complex logic. Moreover, held-out tests may fail to reveal subtle semantic discrepancies between the generated and intended implementations. In addition, when the model produces an incorrect solution, it receives no feedback and has no opportunity to revise or explore alternatives. Finally, existing benchmarks for inductive program synthesis (Gulwani, 2011; Wang et al., 2017; Wong et al., 2021; Deng et al., 2024; Li & Ellis, 2025) are focused on domain-specific tasks and do not assess the ability of LLMs to synthesize functions written in general-purpose programming languages.
@@
To address the limitations of existing evaluation protocols for inductive program synthesis, we introduce CodeARC, the Code Abstraction and Reasoning Challenge, inspired by real-world scenarios such as decompilation and reverse engineering (Lin et al., 2010; Mantovani et al., 2022; Al-Kaswan et al., 2023). In such settings, an agent is given a binary executable (without source code) and must synthesize equivalent source code by observing input-output behavior. Instead of relying on a fixed dataset, the agent can query the binary with new inputs, invoke a differential testing oracle, and use counterexamples for iterative refinement. This setup parallels the classic learnability framework of queries and counterexamples (Angluin, 1987; De la Higuera, 2010), here applied to program synthesis.
Figure˜1 illustrates how CodeARC instantiates this process: LLM-based agents begin with an initial set of input-output examples, query the ground-truth function for more examples, and debug synthesized code using a differential testing oracle. We impose fixed budgets on both the number of observable input-output examples and the number of testing oracle invocations for self-debugging. The task requires agents to proactively generate inputs (function calls) and revise solutions based on feedback (self-reflection). This interactive setup offers a more realistic alternative to prior static evaluation protocols.
We construct the first comprehensive dataset for general-purpose inductive program synthesis, featuring 1114 functions with initial input-output examples. Our benchmark targets general programming tasks and employs two state-of-the-art differential testing tools (Lukasczyk & Fraser, 2022; Etemadi et al., 2024) for correctness evaluation.
@@ 2 Related Work
Most LLM benchmarks, such as HumanEval+ (Chen et al., 2021; Liu et al., 2023b), MBPP+ (Austin et al., 2021; Liu et al., 2023b), APPS (Hendrycks et al., 2021), and others (Li et al., 2022; Liu et al., 2023c; Li et al., 2023; Jain et al., 2024a; Banerjee et al., 2025; Ni et al., 2024b; Liu et al., 2023d; Du et al., 2023; Suresh et al., 2025b; Zhuo et al., 2024; Yin et al., 2022; Lai et al., 2023; Patil et al., 2025; Ugare et al., 2024; 2025; Wei et al., 2025d; Wu et al., 2025), evaluate code generation from natural language. Beyond generation, tasks like I/O prediction (Gu et al., 2024; Liu et al., 2024a), execution prediction (Liu et al., 2023a; La Malfa et al., 2024; Ni et al., 2024a; Ding et al., 2024), bug localization (Suresh et al., 2025a), and program equivalence (Wei et al., 2025a) have also been studied. In contrast, we focus on predicting function bodies purely from input-output examples, without natural language. Prior work (Li & Ellis, 2025; Barke et al., 2024) targets domain-specific tasks, while we introduce a general-purpose benchmark with an interactive evaluation protocol.
@@
LLMs are widely benchmarked on reasoning tasks across domains, including commonsense (Talmor et al., 2018), mathematical (Cobbe et al., 2021), and logical (Han et al., 2022; Miao et al., 2021; Liu et al., 2020; Parmar et al., 2024; Wei et al., 2025e). Inductive reasoning, a core cognitive skill that generalizes from limited examples (Hayes et al., 2010), is increasingly studied in LLMs (Li et al., 2024; Ma et al., 2024; Xiao et al., 2024; Cai et al., 2024; Shao et al., 2024). ARC (Chollet, 2019) is a prominent benchmark for abstract pattern induction. Our work shares this goal but is for inductive program synthesis.
@@
LLM-based agents have shown strong performance in domains like web navigation (Zhou et al., 2024a; b), code generation (Zhang et al., 2023; Jimenez et al., 2024), performance optimization (Wei et al., 2025b; c), and ML experimentation (Huang et al., 2024). They interact with environments, invoke functions, make decisions, and self-reflect (Madaan et al., 2023; Paul et al., 2023; Xie et al., 2023; Huang et al., 2022; Shinn et al., 2023). We introduce the first benchmark to systematically evaluate agents’ capabilities in inductive program synthesis, providing a rigorous testbed for inductive reasoning and program synthesis.
@@ 3.1 Problem Definition of Inductive Program Synthesis
To evaluate whether a synthesized function $\hat{f}$ is correct, we introduce a differential testing oracle $\mathcal{O}$. The oracle takes as input both the synthesized function $\hat{f}$ and the hidden ground-truth function $f^{*}$ and attempts to identify inputs on which their behaviors differ. Formally, the oracle operates as follows:
@@
Unlike fixed held-out test sets used in prior work, the differential testing oracle conditions on both $f^{*}$ and the candidate $\hat{f}$, generating targeted inputs to expose discrepancies. On failure, it returns a counterexample $x\in\mathcal{X}_{\text{test}}$ such that $\hat{f}(x)\neq f^{*}(x)$. Note that program equivalence checking is fundamentally undecidable, and thus no perfect oracle exists. To approximate oracle functionality, we adopt two state-of-the-art differential testing tools, enabling a more robust and practical evaluation than prior work or reliance on a single tool (see Section˜A.3).
@@ 3.2 Interactive Evaluation Protocol for LLM Agents
[…]
@@
[…]
@@ 3.3 Benchmark Preparation
We generate a synthetic dataset for fine-tuning (described in more detail in Section˜3.4) by first collecting 50 seed Python functions that are disjoint from the evaluation set. Using these seeds, we prompt GPT-4o to synthesize a diverse set of new functions, yielding 10,000 candidates. For each generated function, we additionally instruct GPT-4o to produce 10 representative inputs that expose the function’s behavior and highlight patterns in its input-output relationships. These inputs are executed to verify the executability of the functions, and we discard any that fail at this stage. To reduce redundancy, we deduplicate the dataset based on function names, finally resulting in 5,405 unique Python functions used for fine-tuning.
@@ 3.4 Fine-Tuning on Synthetic Data
We evaluate whether fine-tuning on curated synthesis traces, which capture both function calls and reasoning, improves LLM performance on CodeARC, using a distillation approach that imitates the reasoning of a teacher model with access to $f^{*}$, the ground-truth function.
During training, we first run the interactive evaluation protocol described in Section 3.2 with a frozen teacher model. Unlike a standard evaluation, we prepend a set of task-specific instructions $P_{f^{*}}$ to each teacher prompt. This prefix includes the function body of $f^{*}$ and explicitly instructs the teacher to (1) query $f^{*}$ on informative inputs, (2) explain the rationale behind those queries, and (3) synthesize the full implementation of $f^{*}$ only when confident that the correct logic can be inferred from the accumulated input-output pairs. The student model, which is the only model we fine-tune, learns to mimic the teacher’s reasoning and synthesis behavior without seeing the teacher’s prompt that includes the target function.
@@ 4 Experiment Setup
We construct two versions (annotated and anonymized) of the 1114 Python functions, drawn from HumanEval+, MBPP+ (Liu et al., 2023b), and APPS (Hendrycks et al., 2021). Table˜1 summarizes key statistics. Unlike prior work in program synthesis that often focuses on domain specific languages and constrained settings, our benchmark consists of programs written in Python, a general-purpose language that captures a broader range of real-world algorithms and tasks.
For the main evaluation (Section˜5.1), we provide 10 initial input-output examples and set the query budget to 30 input-output pairs and 2 oracle calls ($B_{\text{io}}=30$, $B_{\text{oracle}}=2$), chosen based on practical constraints such as API cost and runtime. Section˜5.3 reports ablation studies on both budgets. We use two state-of-the-art differential testing tools, Pynguin (Lukasczyk & Fraser, 2022) and Mokav (Etemadi et al., 2024). For supervised fine-tuning on synthetic reasoning trajectories, we use gpt-4o as the teacher model and LLaMA-3.1-8B-Instruct as the student. See Appendix˜A for further details, including prompts, fine-tuning parameters, and additional results on the differential testing tools.
@@ 5.1 Main Results
Table˜2 shows the results for 18 large language models on CodeARC. The order is sorted based on the success rate on the anonymized dataset. We also report the average number of observed input-output examples and oracle invocations. Our findings are as follows:
@@
Larger models generally achieve better performance, as shown in Figure˜2 (log-scale x-axis). All model families exhibit scaling trends, though with varying consistency. Llama-3.1 scales steadily, while Llama-3.2 plateaus at larger sizes, likely due to its multimodal focus. Qwen2.5 shows clearer scaling on anonymized data, where reasoning is required over memorization, highlighting model size’s impact on generalization.
@@ 5.2 Do Initial Input-Output Examples Underspecify the Target Function?
To assess whether 10 input-output examples (Li & Ellis, 2025) suffice to specify the target function, we evaluate the first synthesized function of o3-mini, i.e., the strongest model in our study. Functions that pass the initial examples but fail under oracle testing indicate under-specification, motivating the need for additional examples or oracle-guided feedback. As shown in Table˜4, 67.6% pass the initial 10 input-output example tests, but only 38.9% pass the oracle, revealing 640 cases (28.7%) where the initial examples fail to uniquely specify the target function. These findings show that initial input-output examples often under-specify program behavior, motivating additional queries and oracle-guided feedback for reliable evaluation. This motivates the design of our interactive evaluation protocol.
@@ 5.3 Ablation Study: Input-Output Queries and Oracle Feedback
We evaluate o3-mini and o1-mini with input-output budgets of 10, 20, and 30, using the same setup as Section˜5.1. Table˜4 shows that success rates improve consistently with more examples.
@@
We vary the number of allowed oracle invocations and report success rates in Figure˜3 for four models on both datasets. More oracle calls consistently improve performance, showing that counterexamples from differential testing are valuable for guiding iterative refinement.
@@ 5.4 Performance of Fine-Tuned Models
Table 5 shows that fine-tuning the LLaMA-3.1-8B-Instruct model on curated synthesis traces yields consistent improvements across both datasets. The larger gain on the annotated variant suggests that fine-tuning is particularly effective when semantically informative identifiers are present. Notably, this performance gap emerges despite both datasets being evaluated under the same model architecture and training methodology discussed in Section˜3.4. These results indicate that while fine-tuning helps, there remains substantial room for further improvement. This suggests that future research may focus on enhancing the quality and diversity of the fine-tuning dataset, particularly for the anonymized variant, where gains are more limited. Another promising direction is to explore reinforcement learning approaches (Wei et al., 2025c) that optimize for higher synthesis success rates, potentially overcoming limitations of supervised fine-tuning alone.
To assess whether fine-tuning affects general code generation capabilities, we additionally evaluate the models on BigCodeBench (Zhuo et al., 2024). Unlike our inductive program synthesis benchmark, BigCodeBench measures traditional code generation from natural language descriptions, covering diverse function calls and library usage. Our synthetic fine-tuning dataset is constructed independently, with no overlap with BigCodeBench, providing a clean test of the fine-tuned model’s performance on other coding tasks.
On BigCodeBench, the base model achieves a pass@1 score of 40.1%, while the fine-tuned model attains 39.6%, a marginal decrease of 0.5 percentage points. This aligns with prior findings on catastrophic forgetting (Kotha et al., 2024), a well-documented phenomenon in which training on a new task can impair performance on previously learned ones. We consider this small degradation acceptable given the gains in inductive program synthesis.